OpenStudy (dinamix):

Find the inverse transform of each of the followings f(s) = 7/(s+3)^3

OpenStudy (triciaal):

normally to find the inverse switch the x and y then solve for y here f(s) = y and s = x

zepdrix (zepdrix):

Recall a few things: ~Exponential causes a shift ~Transform of t^n is going to be helpful here as well\[\large\rm \mathscr L\{t^n\}=\frac{n!}{s^{n+1}}\]

zepdrix (zepdrix):

If we throw an exponential into the mix,\[\large\rm \mathscr L\{e^{at}t^n\}=\frac{n!}{(s-a)^{n+1}}\]it just shifts s by the amount a,

zepdrix (zepdrix):

\[\large\rm \mathscr L\{e^{\color{orangered}{a}t}\color{royalblue}{t^n}\}=\color{royalblue}{\frac{n!}{(s-\color{orangered}{a})^{n+1}}}\] I'm going to rewrite your expression in a clever way, hopefully you can see what's going on :)\[\large\rm \frac{7}{(s+3)^3}\quad=\quad\frac{7}{2}\cdot\frac{2}{(s-(-3))^{2+1}}\quad=\quad \frac{7}{2}\cdot\color{royalblue}{\frac{2!}{(s-\color{orangered}{(-3)})^{2+1}}}\]

zepdrix (zepdrix):

Hopefully that didn't make it too confusing.

OpenStudy (triciaal):

 how to get this to what you have or is this incorrect?